OpenStudy (anonymous):

the sum of 1st n terms of an A.P is given by Sn=5n^2+3n.Find its nth term.

OpenStudy (amistre64):

Sn = n(a1+an)/2 i think might be useful

OpenStudy (amistre64):

3,6,9,12 6+3=9,+9=18,+12 = 30 4(3+12)/2 = 2(15) = 30

OpenStudy (anonymous):

i know the formula but i don't know what to do n(a1+an)/2=5n^2+3n???

OpenStudy (amistre64):

thats my first instinct :)

OpenStudy (amistre64):

5n^2 + 3n = n (5n+3) 2n(5n+3)/2 n(10n+6)/2

OpenStudy (amistre64):

when n=1, a1=8

OpenStudy (amistre64):

n=2, a1+a2 = 26 8+a2 = 26 a2 = 26-8 = 18

OpenStudy (amistre64):

d = a2-a1 = 18-8 = 10 maybe the terms of the sequence are then: an = 8 + 10(n-1)

OpenStudy (anonymous):

n(a1+an)/2=5n^2+3n n(a1+an)/2=n(5n+3) (a1+an)/2=5n+3 a1+an=10n+6 an=10n+6-a1???

OpenStudy (amistre64):

that looks promising too :)

OpenStudy (amistre64):

if a1 = 8 that leaves us: an = 10n - 2 ; for n>=0

OpenStudy (amistre64):

so if your domain starts at 0, an = 10n - 2 if the domain starts at 1 then, an = 10(n-1) + 8

OpenStudy (amistre64):

same sequence, different domains

OpenStudy (jamesj):

If you have an A.P., with terms a1 = a a2 = a + d a3 = a + 2d an = a + (n-1)d then the sum of the first n terms, Sn = a1 + a2 + ... + an = n.a + n(n-1)d/2 = (d/2)n^2 + (a-d/2)n Therefore, if you're given that Sn = 5n^2 + 3n, it must be that the coefficients of n^2 and n are equal: d/2 = 5 --- (1) a - d/2 = 3 --- (2) From (1) it follows that d = 10 and hence from (2) a - 10/2 = 3 i.e., a = 3 + 5 = 8 Therefore the nth term of the underlying series is an = a + (n-1)d = 8 + 10(n-1)